OpenStudy (anonymous):

Infinite Series - Find the interval of convergence for the following geometric series and, within this interval, find the sum; \[\sum_{n=2}^{Infinity} (x-1)^{n-1}/2^{n+1}\] \[\sum_{n=0}^{Infinity} (x-1)^{n+1}/2^{n+3}\] =(x-1)^n (x-1)^1 / (2^n * 2^3) = (x-1)/8 * (x-1/2)^n Sort of stuck where to go from here. I wasn't sure if A = (x-1)/8 and r= (x-1)/2 and I just use a/(1-r) or what. Any help would be appreciated!

OpenStudy (amistre64):

are you trying to find where the 2 of them meet?

OpenStudy (anonymous):

Not entirely sure what that means.

OpenStudy (anonymous):

The sandwich method? I was using that for Sequences.

OpenStudy (amistre64):

well, ive done intervals of convergence on a single summation; but ive never even considered doing it for 2 at the same time

OpenStudy (amistre64):

i assume it wants the interval where they meet at

OpenStudy (anonymous):

Again, not sure what you mean by 2 at the same time. If you mean the first two equations, the first is N=2 and then I did N = 0 and added two to N.

OpenStudy (anonymous):

If you have another method in mind for solving it I would love to hear it. I think what is getting me is the "X". If it were a 2, instead of X, I could have 1/8 instead of (x-1)/8, which would be my A, and then (x-1)/2 would be my R and I would use a/(1-r)

OpenStudy (amistre64):

\[\lim_{n\to inf}\frac{(x-1)^{n-1}}{2^{n+1}}\frac{2^{n}}{(x-1)^{n-2}}\] \[\lim_{n\to inf}\frac{(x-1)}{2}=\frac{x-1}{2}\]

OpenStudy (amistre64):

i thought you had to sums there to play with; but it looks like you just altered the second one from teh first

OpenStudy (anonymous):

Yeah - random question - how do you get the division to go one on top of the other?

OpenStudy (anonymous):

Thanks for the link, Ill check it out.

OpenStudy (amistre64):

the limit is a ratio of \[\frac{a_n}{a_{n-1}}\] and the fraction thing is the code: \frac{top}{bottom}

OpenStudy (amistre64):

if im reading the site right; the interval of convergence is then: \[\frac{|x-1|}{2}<1\] \[|x-1|<2\] -1 < x < 3